Let {an}



n=1 be a convergent sequence with limit L. Then use the definition of


limit to prove that


lim


n→∞

−an = −L.


Now if L = 0, then use the definition of limit to prove that


limn→∞


(−1)n


an = 0


1
Expert's answer
2021-01-26T04:15:38-0500

Let "\{a_n\}_{n=1}^{\infty}" be a convergent sequence with limit "L". By the definition of limit for any "\epsilon>0" there exists "n_0\in\mathbb N" such that "|a_n-L|<\epsilon" for any "n\ge n_0". Since "|(-a_n)-(-L)|=|-(a_n-L)|=|-1|\cdot|(a_n-L)|=|a_n-L|", we conclude that for any "\epsilon>0" there exists "n_0\in\mathbb N" such that "|(-a_n)-(-L)|=|a_n-L|<\epsilon" for any "n\ge n_0". Therefore, "\lim\limits_{n\to\infty}(-a_n)=-L."


Now if "L=0", then for any "\epsilon>0" there exists "n_0\in\mathbb N" such that "|(-1)^na_n|=|(-1)^n|\cdot|a_n|=|a_n|=|a_n-0|<\epsilon" for any "n\ge n_0". Therefore, by the definition of limit we have that "\lim\limits_{n\to\infty}((-1)^na_n)=0."